# CuClarabel: GPU Acceleration for a Conic Optimization Solver

### v2

@@ Abstract
We present the GPU implementation of the general-purpose interior-point solver Clarabel for convex optimization problems with conic constraints. We introduce a mixed parallel computing strategy that processes linear constraints first, then handles other conic constraints in parallel. This mixed parallel computing strategy currently supports linear, second-order cone, exponential cone, and power cone constraints. We demonstrate that integrating a mixed parallel computing strategy with GPU-based direct linear system solvers enhances the performance of GPU-based conic solvers, surpassing their CPU-based counterparts across a wide range of conic optimization problems. We also show that employing mixed-precision linear system solvers can potentially achieve additional acceleration without compromising solution accuracy.
@@ 1 Introduction
$\mathbb{P}$ =$=\left\{z\ \left|\ A^{T}z=0,~{}z\in\mathcal{K}^{*},~{}\left\langle{b},{z}\right\rangle<0\right.\right\},$ (1)
@@
$\mathbb{D}$ =$=\left\{x\ \left|\ ~{}Px=0,~{}-Ax\in\mathcal{K},~{}\left\langle{q},{x}\right\rangle<0\right.\right\}.$ (2)
Finding an optimal solution with strong duality or detecting a strongly infeasible certificate can be unified into solving a linear complementarity problem with two additional slack variables $\tau,\kappa\geq 0$ [19], which can be reformulated to the following problem [15]:
@@
[…]
@@
[…]
@@ 1.1 Our contribution
We describe a Julia GPU implementation of the general-purpose interior-point solver Clarabel [15]. In our implementation, we support cases where $\mathcal{K}$ is an intersection of atomic cones: zero cones, nonnegative cones, second order cones, exponential cones, and power cones. We propose a mixed parallel computing strategy that parallelizes computing for each type of cone, integrates the CUDSS [21] library for linear system solving, and supports mixed-precision linear system solves for moderate speed improvements. Furthermore, we evaluate our solver against others across a variety of conic optimization problems. Our implementation of CuClarabel is available at https://github.com/cvxgrp/CuClarabel.
@@ 1.2 Related work
GPUs are playing an increasingly significant role in scientific computing. In optimization, GPUs are used to run solvers based on first-order methods. For example, CuPDLP [30] is based on the popular PDHG algorithm [31], which requires only matrix multiplication and addition without the use of direct methods (i.e. it is factorization-free). Solvers that require the solution to linear systems, like SCS [13] and CuOSQP [32], have relied on indirect iterative methods—such as the conjugate gradient (CG), the minimal residual (MINRES) [33], and the generalized minimum residual (GMRES) [34] methods—to solve these systems on GPUs. However, the linear systems solved in first-order methods are generally much better conditioned than those encountered in interior-point methods, where the linear systems become increasingly ill-conditioned as the iterations progress. As an interior-point method approaches higher precision, the number of iterations for each inner indirect linear solves increases significantly, which will eventually offset benefits of GPU parallelism and make GPU-based interior-point methods less preferable compared to CPU-based solvers with direct methods in overall computational time [35]. Recently, NVIDIA released the CUDSS package [36], which provides fast direct methods on GPUs for sparse linear systems. Previous work has integrated CUDSS in a nonlinear optimization solver [37], resulting in significant speed-up on large-scale problems compared to its CPU-based counterpart.
@@ 1.3 Paper outline
In §2, we review and discuss the interior point method used by Clarabel [15] for solving conic optimization problems. In §3, we outline how to implement parallel computation for cone operations and solve linear systems within our GPU solver. §4 details our numerical experiments. We detail the supported cones in §A, how to calculate the barrier functions in §B and scaling matrices for each cone in §C.
2 Overview of Clarabel
We first sketch briefly the main operations used in the Clarabel solver [15] for computing a solution to ($\mathcal{P}$). Solving the equivalent problem ($\mathcal{H}$) amounts to finding a root of the following nonlinear equations
@@ 2 Overview of Clarabel
$(x,z,s,\tau,\kappa)\in\mathcal{F}=\textbf{R}^{n}\times\mathcal{K}\times\mathcal{K}^{*}\times\textbf{R}_{+}\times\textbf{R}_{+}.$
Besides the zero cone that is a linear constraint, other supported conic constraints are smoothed by nonlinear equations in pairs within an interior-point method,
@@
where $\mu$ is the smoothing parameter and $f(\cdot)$ is the logarithmically homogeneous self-concordant barrier (LHSCB) function for cone $\mathcal{K}^{*}$. The barrier functions for different cones are detailed in §B. The trajectory (also called the central path)
[…]
where $v=(x,z,s,\tau,\kappa)$, characterizes the solution of (3) in the right limit $\mu\rightarrow 0$, given an initial point $v^{0}$. After starting from $v^{0}$, the Clarabel solver iterates the following steps for each iteration $k$.
@@
$r_{p}$ $=A\bar{x}-\bar{s}+b,$
$r_{d}$ $=P\bar{x}+A^{\top}\bar{z}+q,$
with corresponding primal and dual objectives
$g_{p}$ $=\frac{1}{2}\bar{x}^{\top}P\bar{x}+q^{T}x,$
$g_{d}$ $=-\frac{1}{2}\bar{x}^{\top}P\bar{x}-b^{T}\bar{z},$
and complementarity slackness $\mu=\frac{{s}^{\top}z+\kappa\tau}{\nu+1}$. The solver returns an approximate optimal point if
$\left\|r_{p}\right\|_{\infty}$ $<\epsilon_{\text{feas}}\max\{1,\left\|b\right\|_{\infty}+\left\|\bar{x}\right\|_{\infty}+\left\|\bar{s}\right\|_{\infty}\}$
$\left\|r_{d}\right\|_{\infty}$ $<\epsilon_{\text{feas}}\max\{1,\left\|q\right\|_{\infty}+\left\|\bar{x}\right\|_{\infty}+\left\|\bar{z}\right\|_{\infty}\}$
$|g_{p}-g_{d}|$ $<\epsilon_{\text{feas}}\max\{1,\min\{|g_{p}|,|g_{d}|\}\}.$
@@
$\left\|A^{T}z\right\|_{\infty}$ $<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty}+\left\|z\right\|_{\infty})(b^{T}z)$
$b^{T}z$ $<-\epsilon_{\text{inf}},$
@@
$\left\|Px\right\|_{\infty}$ $<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty})(b^{T}z)$
$\left\|Ax+s\right\|_{\infty}$ $<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty}+\left\|s\right\|_{\infty})(q^{T}x)$
$q^{T}x$ $<-\epsilon_{\text{inf}}.$
@@
where $H$ is the same scaling matrix as in the CPU version of Clarabel [15] and described in §C. We have shown in [15] that solving (6) reduces to solve the next linear system with two different right-hand sides,
$\underbrace{\begin{bmatrix}P&A^{\top}\\ A&-H\end{bmatrix}}_{K}\left[\begin{array}[]{c|c}\Delta x_{1}&\Delta x_{2}\\ \Delta z_{1}&\Delta z_{2}\end{array}\right]=\left[\begin{array}[]{c|c}d_{x}&-q\\ -(d_{z}-d_{s})&b\end{array}\right].$ (7)
After solving (7), we recover the search direction $\Delta=(\Delta x,\Delta z,\Delta\tau,\Delta s,\Delta\kappa)$ using
$\Delta\tau=\frac{d_{\tau}-d_{\kappa}/\tau+(2P\xi+q)^{\top}\Delta x_{1}+b^{\top}\Delta z_{1}}{\kappa/\tau+\xi^{\top}P\xi-(2P\xi+q)^{\top}\Delta x_{2}-b^{\top}\Delta z_{2}}$
$=\frac{d_{\tau}-d_{\kappa}/\tau+q^{\top}\Delta x_{1}+b^{\top}\Delta z_{1}+2\xi^{\top}P\Delta x_{1}}{\|\Delta x_{2}-\xi\|_{P}^{2}-\|\Delta x_{2}\|_{P}^{2}-q^{\top}\Delta x_{2}-b^{\top}\Delta z_{2}},$ (8a)
@@
$(d_{x},d_{z},d_{\tau})=(1-\sigma)G(x,z,s,\tau,\kappa),\quad d_{\kappa}=\kappa\tau+\Delta\kappa\Delta\tau-\sigma\mu,\\ d_{s}=\left\{\begin{array}[]{cl}W^{\top}\left(\lambda\backslash\left(\lambda\circ\lambda+\eta-\sigma\mu\mathbf{e}\right)\right)&\text{(symmetric)}\\ s+\sigma\mu\nabla f(z)+\eta&\text{(nonsymmetric),}\end{array}\right.\end{aligned}$ (9)
@@
for symmetric cones and the $3$rd-order correction [50]
@@
for nonsymmetric cones. The centering parameter $\sigma$ controls the decreasing speed of both the residual $G(x,z,s,\tau,\kappa)$ and the complementarity slackness $\mu$. which is determined heuristically by value of the affine step size $\alpha_{a}$, which is the maximal value ensuring $v+\alpha_{a}\Delta_{a}\in\mathcal{F}$.
@@
At the end of each iteration $k$, we move the current iterate $v$ along the combined direction $\Delta_{c}$ and obtain the new iterate $v+\alpha_{c}\Delta_{c}$. The combined step size $\alpha_{c}$ should satisfy $v+\alpha_{c}\Delta_{c}\in\mathcal{F}$, and furthermore we have to ensure that the new iterate $v+\alpha_{c}\Delta_{c}$ stays in the neighborhood of the central path (5[50].
@@ 3.1 A primer on GPU programming
Due to their natural parallelism, GPUs differ dramatically from CPUs in the way their transistors are configured. GPUs have a smaller number of caches (blocks where memory access is fast) and instruction processing blocks, and far more, albeit simpler, computational blocks (i.e. arithmetic logic and floating point units). This suggests that CPUs utilize their larger caches to minimize instruction and memory latency within each thread, while GPUs switch between their significantly larger number of threads to hide said latency.
@@
In our approach, we utilize the SIMD structure of the GPU to accelerate our interior-point method. For each iteration of our solving algorithm, we execute the following steps:
  Interior-point method. Main computing steps.
  1.    1.
    Update the variable $v$.
  2.    2.
    Update residuals and objective values for termination check.
  3.    3.
    Update the scaling matrix $H$.
  4.    4.
    Factorize the matrix $K$ in (7).
  5.    5.
    Compute the right-hand sides of (7) for the constant, affine step, and combined steps.
  6.    6.
    Solve the linear systems (7) three times with different right-hand sides.
  7.    7.
    Recover the step direction $\Delta$ via (8).
  8.    8.
    Perform a line search and neighborhood check with respect to the central path.
The matrix factorization (step 4) and the back-solve (step 6) are the most time-consuming parts in each iteration of an interior point method [48] and can be computed using the CUDSS package. Steps 1 and 2 contain matrix addition and multiplication operations that have already been supported in CUDA, while steps 3, 5, 7, and 8 include cone operations that should be tailored for each cone. We will detail how to parallelize each of these steps in the next subsection.

### v3

@@ Abstract
We present the GPU implementation of the general-purpose interior-point solver Clarabel for convex optimization problems with conic constraints. We introduce a mixed parallel computing strategy that processes linear constraints first, then handles other conic constraints in parallel. The GPU solver currently supports linear equality and inequality constraints, second-order cones, exponential cones, power cones and positive semidefinite cones of the same dimensionality. We demonstrate that integrating a mixed parallel computing strategy with GPU-based direct linear system solvers enhances the performance of GPU-based conic solvers, surpassing their CPU-based counterparts across a wide range of conic optimization problems. We also show that employing mixed-precision linear system solvers can potentially achieve additional acceleration without compromising solution accuracy.
@@ 1 Introduction
[…]
@@
[…]
Finding an optimal solution with strong duality or detecting a strongly infeasible certificate can be unified into solving a linear complementarity problem with two additional slack variables $\tau,\kappa\geq 0$ [19], which can be reformulated to the following problem:
@@
with respect to $x,s,z,\tau,\kappa$. It is well-known that when $P=0$, ($\mathcal{H}$) is the homogenous self-dual embedding [20] of ($\mathcal{P}$), and the extension for $P\neq 0$ can be regarded as a homogeneous embedding for linear complementarity problems [21, 22]. Precisely, in our case ($\mathcal{H}$) is homogeneous but not self-dual. In [15] it is also shown that $\mathcal{H}$ is always (asymptotically) feasible, and we can recover either an optimal solution or a strong infeasibility certificate of ($\mathcal{P}$) and $(\mathcal{D})$, depending on the value of the optimal solution $(x^{\star},z^{\star},s^{\star},\tau^{\star},\kappa^{\star})$ to ($\mathcal{H}$):
@@
The interior-point method [23] is a popular choice for solving ($\mathcal{H}$). However, it usually requires to factorize linear systems that are increasingly ill-conditioned. The time complexity of matrix factorization scales with respect to the fill-in that is closely related to nonzeros of the matrix and permutation strategies during factorization. It is thus time-consuming to solve large-scale conic optimization problems with interior-point methods.
@@ 1.1 Our contribution
We contribute a GPU-accelerated implementation of the general-purpose interior-point solver Clarabel [15], in Julia. This implementation supports cases where $\mathcal{K}$ is a product of atomic cones: zero cones, nonnegative cones, second-order cones, exponential cones, power cones and positive semidefinite cones. We propose a mixed parallel computing strategy that parallelizes computing for each type of cone, integrates the cuDSS [24] library for linear system solving, and supports mixed-precision linear system solves for moderate speed improvements. Furthermore, we evaluate our solver against others across a variety of conic optimization problems. Our implementation of CuClarabel is open-sourced on GitHub [^1] https://github.com/oxfordcontrol/Clarabel.jl/tree/CuClarabel and can be easily accessed through CVXPY [25].
@@ 1.2 Related work
GPUs are playing an increasingly significant role in scientific computing. In optimization, GPUs are used to run solvers based on first-order methods. For example, CuPDLP [34] is based on the popular PDHG algorithm [35], which requires only matrix multiplication and addition without the use of direct methods (i.e., it is factorization-free). Solvers that require the solution to linear systems, like SCS [13] and CuOSQP [36], have relied on indirect iterative methods—such as the conjugate gradient (CG), the minimal residual (MINRES) [37], and the generalized minimum residual (GMRES) [38] methods—to solve these systems on GPUs. However, the linear systems solved in first-order methods are generally much better conditioned than those encountered in interior-point methods, where the linear systems become increasingly ill-conditioned as the iterations progress. As an interior-point method approaches higher precision, the number of iterations for each inner indirect linear solves increases significantly, which will eventually offset benefits of GPU parallelism and make GPU-based interior-point methods less preferable compared to CPU-based solvers with direct methods in overall computational time [39]. Recently, NVIDIA released the cuDSS package [40], which provides fast direct methods on GPUs for sparse linear systems. Previous work has integrated cuDSS in a nonlinear optimization solver, MADNLP [41], resulting in significant speed-up on large-scale problems compared to its CPU-based counterpart.
@@ 1.3 Paper outline
In §2, we review supported cones and discuss the interior point method used by Clarabel [15] for solving conic optimization problems. In §3, we outline how to implement parallel computation for cone operations and solve linear systems within our GPU solver. §4 details our numerical experiments. We detail the scaling matrices for each cone in §A.
2 Backgrounds
We first review supported cones in CuClarabel with the corresponding barrier functions, along with their dual cones. We then briefly sketch the main operations used by the Clarabel solver [15] to compute a solution to ($\mathcal{P}$).
2.1 Supported cones
We support the following atomic cones in our GPU solver:
  • The zero cone, defined as
    The dual cone of the zero cone is $\left(\{0\}^{n}\right)^{*}=\textbf{R}^{n}$.
  • The nonnegative cone, defined as
  • The positive semidefinite cone $\mathcal{K}_{\mathrm{psd}}^{n}$ is defined as
    $\mathcal{K}_{\mathrm{psd}}^{n}=\left\{X\in\mathbb{S}^{n}\ \middle|\ X\succeq 0\right\},$
    where $\mathbb{S}^{n}$ denotes the space of $n\times n$ symmetric matrices, and $X\succeq 0$ indicates that $X$ is positive semidefinite. The positive semidefinite cone is self-dual, i.e., $\mathcal{K}_{\mathrm{psd}}^{n}=\left(\mathcal{K}_{\mathrm{psd}}^{n}\right)^{*}$. [^2] In the current solver implementation, all positive semidefinite (PSD) cones are restricted to have dimension at most $32$, and all PSD cones within a given problem must share the same dimension.
  • The exponential cone, a $3$-dimensional cone defined as
    with its dual cone given by
    with its dual cone given by
2.2 Interior point method
Solving the problem ($\mathcal{H}$) amounts to finding a root of the following nonlinear equations
@@ 2.2 Interior point method
$(x,z,s,\tau,\kappa)\in\mathcal{F},$
where $\mathcal{F}:=\textbf{R}^{n}\times\mathcal{K}^{*}\times\mathcal{K}\times\textbf{R}_{+}\times\textbf{R}_{+}$ defines the region of cone constraints. Besides the zero cone that is a linear constraint, other supported conic constraints are smoothed by nonlinear equations in pairs within an interior-point method,
@@
where $\mu$ is the smoothing parameter and $f(\cdot)$ is the logarithmically homogeneous self-concordant barrier (LHSCB) function for cone $\mathcal{K}^{*}$. The barrier functions for different cones are:
  1. 1.
    Nonnegative cone $\textbf{R}_{+}^{n}$ of degree $n$:
    $f(z)=-\sum_{i\in\llbracket n\rrbracket}\log(z_{i}),\ z\in\textbf{R}_{+}^{n}.$ (5)
  2. 2.
    Second order cone $\mathcal{K}_{q}^{n}$ of degree $1$:
    $f(z)=-\frac{1}{2}\log\left(z_{1}^{2}-\sum_{i=2}^{n}z_{i}^{2}\right),\ z\in\mathcal{K}_{q}^{n}.$ (6)
  3. 3.
    Positive semidefinite cone $\mathcal{K}_{\mathrm{psd}}^{n}$ of degree $n$:
    $f(X)=-\log\det(X),\quad X\in\mathcal{K}_{\mathrm{psd}}^{n}.$ (7)
  4. 4.
    Dual exponential cone $\mathcal{K}_{\mathrm{exp}}^{*}$ of degree $3$:
    $f(z)=-\log\left(z_{2}-z_{1}-z_{1}\log\left(\frac{z_{3}}{-z_{1}}\right)\right)-\log(-z_{1})-\log(z_{3}),\ z\in\mathcal{K}_{\mathrm{exp}}^{*}.$ (8)
  5. 5.
    Dual power cone $\mathcal{K}_{\mathrm{pow}}^{*}$ of degree $3$:
    $f(z)=-\log\left(\left(\frac{z_{1}}{\alpha}\right)^{2\alpha}\left(\frac{z_{2}}{1-\alpha}\right)^{2(1-\alpha)}-z_{3}^{2}\right)-(1-\alpha)\log(z_{1})-\alpha\log(z_{2}),\ z\in\mathcal{K}_{\mathrm{pow}}^{*}.$ (9)
The trajectory (also called the central path)
$\begin{aligned} G(v)=\mu G(v^{0}),\\ s=-\mu\nabla f(z),\quad\tau\kappa=\mu,\end{aligned}$ (10)
where $v=(x,z,s,\tau,\kappa)$, characterizes the solution of (3) in the right limit $\mu\rightarrow 0^{+}$, given an initial point $v^{0}$. After starting from $v^{0}$, the Clarabel solver iterates the following steps for each iteration $k$:
@@
$\begin{aligned} r_{p}&=-A\bar{x}-\bar{s}+b,\\ r_{d}&=P\bar{x}+A^{T}\bar{z}+q,\\ g_{p}&=\frac{1}{2}\bar{x}^{T}P\bar{x}+q^{T}\bar{x},\\ g_{d}&=-\frac{1}{2}\bar{x}^{T}P\bar{x}-b^{T}\bar{z}\end{aligned}$ (11)
and complementarity slackness $\mu=\frac{{s}^{T}z+\kappa\tau}{\nu+1}$, where $\nu$ is the degree of cone $\mathcal{K}$. The solver returns an approximate optimal point if
$\begin{aligned} \left\|r_{p}\right\|_{\infty}&<\epsilon_{\text{feas}}\max\{1,\left\|b\right\|_{\infty}+\left\|\bar{x}\right\|_{\infty}+\left\|\bar{s}\right\|_{\infty}\}\\ \left\|r_{d}\right\|_{\infty}&<\epsilon_{\text{feas}}\max\{1,\left\|q\right\|_{\infty}+\left\|\bar{x}\right\|_{\infty}+\left\|\bar{z}\right\|_{\infty}\}\\ |g_{p}-g_{d}|&<\epsilon_{\text{feas}}\max\{1,\min\{|g_{p}|,|g_{d}|\}\}.\end{aligned}$ (12)
@@
[…]
@@
$\begin{aligned} \left\|Px\right\|_{\infty}&<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty})(b^{T}z)\\ \left\|Ax+s\right\|_{\infty}&<-\epsilon_{\text{inf}}\max(1,\left\|x\right\|_{\infty}+\left\|s\right\|_{\infty})(q^{T}x)\\ q^{T}x&<-\epsilon_{\text{inf}}.\end{aligned}$ (13b)
@@
where $\xi=x\tau^{-1}$, and $H$ is the same positive-definite scaling matrix as in the CPU version of Clarabel [15], also described in §A. We have shown in [15] that solving (14) reduces to solve the next linear system with two different right-hand sides,
$\underbrace{\begin{bmatrix}P&A^{T}\\ A&-H\end{bmatrix}}_{K}\left[\begin{array}[]{c|c}\Delta x_{1}&\Delta x_{2}\\ \Delta z_{1}&\Delta z_{2}\end{array}\right]=\left[\begin{array}[]{c|c}d_{x}&-q\\ -(d_{z}-d_{s})&b\end{array}\right].$ (15)
Since both $P,H$ are positive semidefinite, adding a positive regularization to them makes $K$ symmetric quasi-definite and thus strongly factorizable via the $LDL^{T}$ decomposition [51]. After solving (15), we recover the search direction $\Delta=(\Delta x,\Delta z,\Delta\tau,\Delta s,\Delta\kappa)$ using
$\Delta\tau=\frac{d_{\tau}-d_{\kappa}/\tau+(2P\xi+q)^{T}\Delta x_{1}+b^{T}\Delta z_{1}}{\kappa/\tau+\xi^{T}P\xi-(2P\xi+q)^{T}\Delta x_{2}-b^{T}\Delta z_{2}}$
$=\frac{d_{\tau}-d_{\kappa}/\tau+q^{T}\Delta x_{1}+b^{T}\Delta z_{1}+2\xi^{T}P\Delta x_{1}}{\|\Delta x_{2}-\xi\|_{P}^{2}-\|\Delta x_{2}\|_{P}^{2}-q^{T}\Delta x_{2}-b^{T}\Delta z_{2}},$ (16a)
@@
in the combined step (predictor+corrector), where $\lambda=W^{-T}s=Wz$ and $\mathbf{e}$ is the idempotent for a symmetric cone with the product operator $\circ$’ and its inverse operator $\backslash$’ [52]. Here $\eta$ denotes a higher-order correction term, which is a heuristic technique that can significantly accelerate the convergence of interior-point methods [53]. We set it to the Mehrotra correction [54]
@@
for symmetric cones and the 33rd-order correction [55]
@@
for nonsymmetric cones. The centering parameter $\sigma$ controls the rate at which both the residual $G(x,z,s,\tau,\kappa)$ and the complementarity measure $\mu$ decrease. It is determined heuristically based on the value of the affine step size $\alpha_{a}$, defined as the maximal step size ensuring that $v+\alpha_{a}\Delta_{a}\in\mathcal{F}$. Computing $\alpha_{a}$ is non-trivial in the presence of exponential and power cones. As a practical alternative, we perform a backtracking line search with a shrinking ratio of $0.8$ to determine a feasible $\alpha_{a}$ for power and exponential cones.
@@
At the end of each iteration $k$, we move the current iterate $v$ along the combined direction $\Delta_{c}:=(x,z,s,\tau,\kappa)$ and obtain the new iterate $v+\alpha_{c}\Delta_{c}$. As discussed, the combined step size $\alpha_{c}$ should satisfy $v+\alpha_{c}\Delta_{c}\in\mathcal{F}$. Furthermore, we must ensure that the new iterate $v+\alpha_{c}\Delta_{c}$ stays in the neighborhood $\mathcal{N}(\beta)$ of the central path (10[55],
$\mathcal{N}(\beta)=\left\{(s,z)\in\mathcal{K}\times\mathcal{K}^{*}\;\middle|\;\nu_{i}\left\langle\nabla f^{*}(s_{i}),\nabla f(z_{i})\right\rangle^{-1}\geq\beta\mu,\;i=1,\ldots,p\right\},$ (20)
where $\beta$ is set to $10^{-6}$.
@@ 3.1 A primer on GPU programming
Due to their natural parallelism, GPUs differ dramatically from CPUs in the way their transistors are configured. GPUs have a smaller number of caches (blocks where memory access is fast) and instruction processing blocks, and far more, albeit simpler, computational blocks (i.e., arithmetic logic and floating point units). This suggests that CPUs utilize their larger caches to minimize instruction and memory latency within each thread, while GPUs switch between their significantly larger number of threads to hide said latency.
@@
3.2 GPU implementation for interior point method
We detail our GPU implementation of the primal-dual interior point method in Algorithm 1, which is divided into two phases: setup and solve. The setup phase involves data equilibration and solver initialization, while the solve phase implements the classical primal-dual interior point method. Synchronization is required at the end of each step in Algorithm 1.
Algorithm 1 Primal-dual interior-point method
0: Parameter inputs $P,A,q,b$ and cone information $\mathcal{K}$. //Setup phase
1: Equilibrate $P,A,q,b$.
2: Initialize solver structure $s$, including matrix $K$ and cuDSS solver.//Solve phase
3:while termination check (12) or (13) is not satisfied do
4:  Update the scaling matrix $H$ as in Appendix A.
5:  Factorize the matrix $K$ in (15).
6:  Compute the right-hand residual $d$ (17) for the affine step and solve (15) to obtain $\Delta_{a}$ via (16).
7:  Compute the affine step $\alpha_{a}$ satisfying $v+\alpha_{a}\Delta_{a}\in\mathcal{F}$.
8:  Compute the centering parameter $\sigma=(1-\alpha_{a})^{3}$.
9:  Compute the right-hand residual $d$ (18) for the combined step and solve (15) to obtain $\Delta_{c}$ via (16).
10:  Compute the combined step $\alpha_{c}$ satisfying $v+\alpha_{c}\Delta_{c}\in\mathcal{F}$ with the neighborhood check.
11:  Update the variable $v\leftarrow v+0.99\alpha_{c}\Delta_{c}$.
12:  Update residuals and objective values in (11).
13:end while
The matrix factorization (step 5) and the back-solve (steps 6 and 9) are the most time-consuming parts in each iteration of an interior point method, and can be computed using the cuDSS package. Variable and information update (steps 11 and 12) contain matrix addition and multiplication operations that have already been supported in CUDA. Steps 4, 6, 7, 9, and 10 include cone operations that should be tailored for each cone, but we can abstract them into the same framework for parallelism. We will detail how to parallelize each of these steps under a unified framework in the next subsection.